# Oblique boundary value problems for degenerate Hessian quotient type equations

### v1

@@ 3. Gradient estimates
Let $u\in C^{3}(\Omega)\cap C^{2}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a smooth bounded domain $\Omega\subset\mathbb{R}^{n}$. Assume $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$ and $\sup_{\Omega}|u|\leq M$, where $M$ is a positive constant. If $G(x,z,p)$ is the form of (1.4), suppose that $\beta$ is a smooth unit vector field along $\partial\Omega$ with $\beta,\nu\right>\geq\beta_{0}>0$ for some positive constant $\beta_{0}$. Then, for any $0\leq l<k<n$,
@@
where $C>0$ is a constant depending only on $M,n,k,l,\gamma,\overline{\Omega},\beta_{0},|\beta|_{C^{3}},|f^{\frac{1}{k-l}}|_{C^{1}},|\varphi|_{C^{3}}.$
@@
Due to Dong-Wei [7], we have already known interior gradient estimates, so we only need to obtain gradient estimates near the boundary. Denote $\Omega_{\mu}=\left\{x\in\Omega\;|\;d(x)\leq\mu<\mu_{0}\leq 1\right\}$, here $\mu,\;\mu_{0}$ are positive constants to be determined later. Then, extend $\beta$ to $\Omega_{\mu}$ smoothly such that $\beta,Dd\right>=\cos\theta\geq\beta_{0}>0$.
Take a unit normal moving frame along $\partial\Omega$, denoted by $\{e_{1},e_{2},\cdots,e_{n-1},\nu\}$, then
@@
where $\beta_{\nu}=\left<\beta,\nu\right>=\cos\theta\geq\beta_{0}>0$. Denote by $d(x)=dist(x,\partial\Omega)$ the distance from $x$ to $\partial\Omega$ and let $w(x)=u(x)-\frac{\varphi(x,z)d(x)}{\cos\theta}$. On $\partial\Omega$, since $\varphi(x,z)=u_{\nu}\beta_{\nu}+\sum_{q=1}^{n-1}u_{q}\beta_{q}$, we have
(3.3) $w_{\nu}=-\sum_{q=1}^{n-1}w_{q}\frac{\beta_{q}}{\beta_{\nu}}.$
Applying the fact $|\beta|^{2}=1$ and $\cos\theta>0$,
(3.4) $w_{\nu}^{2}\leq|Dw|^{2}\sin^{2}\theta.$
Consider an auxiliary function
@@
where
(3.6) $\psi(x)=|Dw|^{2}-\left(\sum_{i=1}^{n}w_{i}d_{i}\right)^{2}=\sum_{i,j=1}^{n}\left(\delta_{ij}-d_{i}d_{j}\right)w_{i}w_{j}=\sum_{i,j=1}^{n}C^{ij}w_{i}w_{j}>0,$
(3.7) $g(u)=\frac{1}{4}\ln\frac{1}{(3M-u)},$
and $\alpha_{1}$ is a large enough positive constant to be determined later. Suppose that $\widetilde{G}(x)$ attains the maximum at $x_{0}\in\overline{\Omega}_{\mu}$, according to interior gradient estimates, we can only consider the following two cases.
$\mathbf{Case\;\mathrm{\mathbf{I}}}$: $x_{0}\in\partial\Omega$. Choosing a coordinate around $x_{0}$ such that $\nu=\frac{\partial}{\partial x_{n}}$ and $\frac{\partial}{\partial x_{s}}\;(s=1,2,\cdots,n-1)$ are tangent to $\partial\Omega$, then
@@
here $\kappa_{s}$ are principal curvatures of $\partial\Omega$ at $x_{0}$. A direct calculation yields
(3.8) $w_{\nu i}=w_{i\nu}+\sum_{j=1}^{n}\mathcal{K}_{ij}w_{j},\quad\text{for}\;i=1,2,\cdots,n,$
where $\mathcal{K}_{ij}$ denoted by the Weingarten matrix of the boundary with respect to $\nu$. In order to derive the later calculations, we deal with $\psi_{\nu}$ in advance
(3.9) $\psi_{\nu}(x_{0})=\psi_{n}=-\frac{2\sum_{s,q=1}^{n-1}w_{s}w_{sq}\beta_{q}}{\beta_{\nu}}-2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{s}-2\sum_{s,q=1}^{n-1}\mathcal{K}_{sq}w_{s}w_{q}.$
According to the fact that $x_{0}$ is the maximum point of $\widetilde{G}(x)$ on the boundary, we have
(3.10) $0=\widetilde{G}_{s}(x_{0})=\frac{\psi_{s}}{\psi}+g^{\prime}u_{s}+\alpha_{0}d_{s}=\frac{\psi_{s}}{\psi}+g^{\prime}u_{s},\quad\text{for}\;s=1,2,\cdots,n-1,$
and
(3.11) $0\geq\widetilde{G}_{\nu}(x_{0})=\widetilde{G}_{n}=\frac{\psi_{n}}{\psi}+g^{\prime}u_{n}+\alpha_{0}.$
By (3.9) and (3.10), we obtain
(3.12) $0\geq\widetilde{G}_{n}=$ $\;-\frac{2w_{n}\sum_{s,q=1}^{n-1}w_{s}\beta_{q}d_{sq}}{\beta_{\nu}\psi}-\frac{2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{p}}{\psi}$
$\;-\frac{2\sum_{s,q=1}^{n-1}\mathcal{K}_{sq}w_{s}w_{q}}{\psi}+\frac{g^{\prime}\varphi}{\cos\theta}+\alpha_{0}$
$=$ $\;-\frac{2\sum_{i,k=1}^{n-1}\kappa_{i}w_{i}w_{k}\beta_{i}\beta_{k}}{\beta_{\nu}^{2}\sum_{i,k=1}^{n}C^{ik}w_{i}w_{k}}-\frac{2\sum_{i,k=1}^{n-1}w_{i}w_{k}\left(\frac{\beta_{k}}{\beta_{\nu}}\right)_{i}}{\sum_{i,k=1}^{n}C^{ik}w_{i}w_{k}}$
$\;-\frac{2\sum_{s,q=1}^{n-1}\mathcal{K}_{sq}w_{s}w_{q}}{\sum_{i,k=1}^{n}C^{ik}w_{i}w_{k}}+\frac{g^{\prime}\varphi}{\cos\theta}+\alpha_{0}$
$\geq$ $\;-C+\alpha_{0}>0,$
where the last inequality comes from taking $\alpha_{0}$ large enough depending on $|\beta|_{C^{1}}$, $n$, $\beta_{0}$ and the geometry of $\partial\Omega$. Thus this case does not occur at all.
$\mathbf{Case\;\mathrm{\mathbf{II}}}$: $x_{0}\in\Omega_{\mu}$. Assume that $|Du|$ attains its boundary maximum at $y_{0}\in\partial\Omega$ and
@@
Since $\widetilde{G}(x_{0})\geq\widetilde{G}(y_{0})$,
(3.14) $\psi(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{0})\text{e}^{-\frac{1}{4}}}{4}|Du|^{2}(y_{0}).$
Then
(3.15) $\psi(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{0})}{8C\text{e}}|Du|^{2}(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{0})}{16C\text{e}}|Dw|^{2}(x_{0}):=C_{0}|Dw|^{2}(x_{0}),$
where denotes $C_{0}\in(0,1)$. Choosing the coordinate such that matrix $D^{2}u(x_{0})$ is diagonal. For $k=1,2,\cdots,n$, denote $T_{k}=\sum_{l=1}^{n}C^{kl}w_{l}$ and $\overrightarrow{T}=(T_{1},T_{2},\cdots,T_{n})$, it is direct to observe that $\left|\overrightarrow{T}\right|\leq|Dw|$ and $\psi\leq\left|\overrightarrow{T}\right||Dw|$. Then
@@
Without loss of generality, we further assume that $T_{1}w_{1}\geq\frac{C_{0}}{n}|Dw|^{2}>0$ and therefore we have $\frac{w_{1}}{T_{1}}\geq\frac{C_{0}}{n}$. Let $\mu$ is small such that
@@
denote ${C^{kl}}_{,i}=\frac{\partial C^{kl}}{\partial x_{i}}$ and ${C^{kl}}_{,ij}=\frac{\partial^{2}C^{kl}}{\partial x_{i}\partial x_{j}}$, especially for $i=1$ in equation (3.18), we obtain
(3.19) $\sum_{l=1}^{n}T_{l}w_{l1}=-\frac{\psi}{2}(g^{\prime}u_{1}+\alpha_{1}d_{1})-\sum_{k,l=1}^{n}\frac{{C^{kl}}_{,1}}{2}w_{k}w_{l}.$
Take $\mu$ small enough and $|Dw|$ large enough, then
(3.20) $u_{11}\left(1-\varphi_{z}\frac{d}{\cos\theta}\right)<0.$
Now, it is turn to deal with second order derivatives of $\widetilde{G}$. Using the fact that $F$ is homogeneous of degree $(k-l)$, we have
(3.21) $0\geq F^{ij}\widetilde{G}_{ij}(x_{0})=&\;\frac{F^{ij}\left(\sum_{k,l=1}^{n}C^{kl}w_{k}w_{l}\right)_{ij}}{\psi}-2\alpha_{1}g^{\prime}F^{ij}u_{i}d_{j}-\alpha_{1}^{2}F^{ij}d_{i}d_{j}\\ &\;+\left[g^{\prime\prime}-\left(g^{\prime}\right)^{2}\right]F^{ij}u_{i}u_{j}+g^{\prime}F^{ij}u_{ij}+\alpha_{1}F^{ij}d_{ij}\\ \geq&\;\frac{F^{ij}\left(\sum_{k,l=1}^{n}C^{kl}w_{k}w_{l}\right)_{ij}}{\psi}-2\alpha_{1}^{2}F^{ij}d_{i}d_{j}+\alpha_{1}F^{ij}d_{ij}\\ &\;+\left[g^{\prime\prime}-2\left(g^{\prime}\right)^{2}\right]F^{ij}u_{i}u_{j}+g^{\prime}(k-l)f\\ :=&\;\text{I}+\text{II}+\text{III}+\text{IV}+\text{V}.\end{aligned}$
@@
where $K_{0}$ is a positive constant related to the geometry of $\partial\Omega$.
Next, we just need to deal with the term I
@@
For the term $I_{2}$, by (2.4), (3.16) and Lemma 2.5, we derive
@@
$\;-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|.$
Similarly,
@@
(3.29) $\mathcal{H}=2\sum_{i=1}^{n}F^{ii}C^{ii}u_{ii}^{2}-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|,$
and we will bound $\mathcal{H}$ from below in the following.
Without loss of generality, we can set $C^{11}$ be the smallest of $\left\{C^{ii}\right\}_{i=1}^{n}$. Then, we have $C^{ii}\geq\frac{1}{2}$ for any $i\geq 2$. otherwise it follows that $\sum_{i=1}^{n}C^{ii}<n-1$, which contradicts with $\sum_{i=1}^{n}C^{ii}=n-1$. A simple calculation yields
(3.30) $\mathcal{H}\geq-C\sum_{i=1}^{n}F^{ii}\left(d|Dw|^{2}+|Dw|\right)^{2}.$
Plugging (3.30) into (3.28), then
@@
Substituting (3.22) and (3.31) into (3.21), we arrive at
(3.32) $0\geq\left\{C_{2}\left[g^{\prime\prime}-2\left(g^{\prime}\right)^{2}\right]|Dw|^{2}-C_{3}\mu|Dw|^{2}-C|Dw|\right\}\sum_{i=1}^{n}F^{ii}.$
Let $\mu$ be small enough such that $C_{2}\left(\frac{1}{16M}\right)^{2}|Dw|^{2}-C|Dw|\leq 0$, then the proof of Theorem 3.1 have been finished.∎

### v2

@@ 3. $C^{0}$ and $C^{1}$ estimates
Assume $\varphi$ is a smooth function and $f_{u}\geq 0$, $\varphi_{u}\geq\gamma_{0}>0$. If there exists a subsolution $\underline{u}\in C^{2}(\overline{\Omega})$ satisfying problem (1.7) and an admissible solution $u\in C^{2}(\overline{\Omega})$ satisfying problem (1.8) in a bounded $C^{2}$ domain $\Omega\subset\mathbb{R}^{n}$, then
$\max_{\overline{\Omega}}|u|\leq C,$
where $C$ depends on $\gamma_{0}$, $|\underline{u}|_{C^{0}}$, $|\varphi|_{C^{1}}$.
Proof.
From the $(k-l)$-degree homogeneity of the operator $F$, we derive $\sum_{i,j}F^{ij}u_{ij}\geq 0$. Consequently, the maximum of $u$ is attained at a boundary point $x_{0}\in\partial\Omega$ with $u(x_{0})>0$, otherwise, the upper bound would be obtained. This implies
$0\geq D_{\beta}u(x_{0})$ $=\varphi(x_{0},u(x_{0}))-\varphi(x_{0},0)+\varphi(x_{0},0)$
$=\varphi_{u}(x_{0},tu(x_{0}))u(x_{0})+\varphi(x_{0},0)$
$\geq\gamma_{0}u(x_{0})+\varphi(x_{0},0),$
where $10<t<$ 1. Then $u(x_{0})\leq\frac{-\varphi(x_{0},0)}{\gamma_{0}}\leq C$.
Under the condition $f_{u}\geq 0$, the maximum principle yields that $\underline{u}-u$ achieves its maximum at a boundary point $x_{1}\in\partial\Omega$, thus
$0\geq D_{\beta}(\underline{u}-u)(x_{1})$ $=\varphi(x_{1},\underline{u}(x_{1}))-\varphi(x_{1},u(x_{1}))$
$=\varphi_{u}[x_{1},tu(x_{1})+(1-t)\underline{u}(x_{1})](\underline{u}-u)(x_{1}),$
where $10<t<$ 1. Since $\varphi_{u}>0$, for any $x\in\overline{\Omega}$
$(\underline{u}-u)(x)\leq(\underline{u}-u)(x_{1})\leq 0.$
Hence we finish the proof. ∎
Theorem 3.2.
Let $u\in C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.8) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Assume that $0\leq f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$, the oblique boundary value condition $u_{\beta}=\varphi(x,u)$ satisfies $\beta,\nu\right>\geq\beta_{0}>0$, where $\beta$ is a smooth unit vector field along $\partial\Omega$, $\nu$ is the unit inner normal vector field along $\partial\Omega$, $\varphi$ is a smooth function defined on $\overline{\Omega}\times\mathbb{R}$. Then for $\gamma\geq 1$, $0\leq l<k<n$
@@
where $C$ is a positive constant depending only on $n,k,l,\gamma,\overline{\Omega},\beta_{0},|\beta|_{C^{3}},|f^{\frac{1}{k-l}}|_{C^{1}},|\varphi|_{C^{3}}$.
@@
Building upon the interior gradient estimates established by Dong-Wei [7], we focus our analysis on deriving gradient estimates near the boundary. Denote $\Omega_{\mu}=\left\{x\in\Omega\;|\;d(x)=dist(x,\partial\Omega)\leq\mu\right\}$, here $\mu$ is a small positive constants to be determined later. Now extend $\beta$ to $\Omega_{\mu}$ smoothly such that $\beta,Dd\right>=\cos\theta\geq\beta_{0}>0$.
Taking a unit normal moving frame along $\partial\Omega$ denoted by $\{e_{1},e_{2},\cdots,e_{n-1},\nu\}$, the oblique vector field decomposes as
@@
where $\beta_{\nu}=\left<\beta,\nu\right>=\cos\theta\geq\beta_{0}>0$, thus $\varphi(x,u)=u_{\nu}\beta_{\nu}+\sum_{q=1}^{n-1}u_{q}\beta_{q}$ on $\partial\Omega$. Defining $w(x)=u(x)-\frac{\varphi(x,u)d(x)}{\cos\theta}$, on $\partial\Omega$ we get
(3.3) $w_{\nu}=-\sum_{q=1}^{n-1}w_{q}\frac{\beta_{q}}{\beta_{\nu}},$
(3.4) $w_{\nu}^{2}\leq\sin^{2}\theta|Dw|^{2}.$
Set an auxiliary function
@@
where $\alpha_{1}$ is a sufficiently large positive constant and
(3.6) $\psi(x)=|Dw|^{2}-\left(\sum_{i=1}^{n}w_{i}d_{i}\right)^{2}=\sum_{i,j=1}^{n}\left(\delta_{ij}-d_{i}d_{j}\right)w_{i}w_{j}=\sum_{i,j=1}^{n}C^{ij}w_{i}w_{j},$
(3.7) $g(u)=\frac{1}{4}\ln\frac{1}{(3M-u)},\quad M=\max_{\overline{\Omega}}u,$
then we derive
$\frac{1}{16M}\leq g^{\prime}\leq\frac{1}{8M},\quad g^{\prime\prime}-2(g^{\prime})^{2}\geq\frac{1}{128M^{2}}.$
Suppose $\widetilde{G}(x)$ achieves its maximum value at $x_{0}\in\overline{\Omega}_{\mu}$. In view of the interior gradient estimates, the analysis reduces to the following two cases.
$\mathbf{Case\;\mathrm{\mathbf{I}}}$: $x_{0}\in\partial\Omega$. We choose a coordinate around $x_{0}$ such that $\nu=\frac{\partial}{\partial x_{n}}$ and $\frac{\partial}{\partial x_{s}}$ are tangent to $\partial\Omega$, then
@@
here $\kappa_{s}$ are principal curvatures of $\partial\Omega$ at $x_{0}$. By virtue of (3.4)
$\psi=|Dw|^{2}-(w_{n}d_{n})^{2}\geq\cos^{2}\theta|Dw|^{2}\geq\beta_{0}^{2}|Dw|^{2}>0.$
A direct calculation yields
(3.8) $w_{\nu s}=w_{s\nu}+\sum_{i=1}^{n}\mathcal{K}_{is}w_{i},$
where $\mathcal{K}_{is}$ denoted by the Weingarten matrix of the boundary with respect to $\nu$. We firstdeal with $\psi_{\nu}$ for later computations
(3.9) $\psi_{\nu}(x_{0})=\psi_{n}=-2\sum_{s,q=1}^{n-1}w_{s}w_{sq}\frac{\beta_{q}}{\beta_{\nu}}-2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{s}-2\sum_{s=1}^{n-1}\sum_{i=1}^{n}\mathcal{K}_{si}w_{s}w_{i}.$
At $x_{0}$
(3.10) 00 0 $=\widetilde{G}_{s}=\frac{\psi_{s}}{\psi}+g^{\prime}u_{s}+\alpha_{1}d_{s}=\frac{\psi_{s}}{\psi}+g^{\prime}u_{s},$
(3.11) 00 0 $\geq\widetilde{G}_{n}=\frac{\psi_{n}}{\psi}+g^{\prime}u_{n}+\alpha_{1}.$
By (3.9)-(3.11)
(3.12) $0\geq$ $\;-\frac{2w_{n}\sum_{s=1}^{n-1}\sum_{i=1}^{n}w_{i}\beta_{s}d_{is}}{\beta_{n}\psi}-\frac{2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{s}}{\psi}$
$\;-\frac{2\sum_{s=1}^{n-1}\sum_{i=1}^{n}\mathcal{K}_{si}w_{s}w_{i}}{\psi}+\frac{g^{\prime}\varphi}{\cos\theta}+\alpha_{1}$
$=$ $\,-\frac{2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{s}}{\psi}-\frac{2\sum_{s=1}^{n-1}\sum_{i=1}^{n}\mathcal{K}_{si}w_{s}w_{i}}{\psi}$
$\,+\frac{2\sum_{s=1}^{n-1}w_{s}\beta_{s}^{2}\kappa_{s}^{2}}{\beta_{n}^{2}\psi}+\frac{g^{\prime}\varphi}{\cos\theta}+\alpha_{1}$
$\geq$ $\;-C+\alpha_{1},$
we immediately arrive at a contradiction from taking $\alpha_{1}$ large enough.
$\mathbf{Case\;\mathrm{\mathbf{II}}}$: $x_{0}\in\Omega_{\mu}$. Recalling $w=u-\frac{\varphi d}{\cos\theta}$, direct calculation yields
(3.13) $w_{k}=$ $\left(1-\frac{\varphi_{z}d}{\cos\theta}\right)u_{k}+O(1),$
$w_{ki}=$ $\left(1-\frac{\varphi_{z}d}{\cos\theta}\right)u_{ki}+O(d|Dw|^{2}+|Dw|+1),$
$w_{lij}=$ $\left(1-\frac{\varphi_{z}d}{\cos\theta}\right)u_{lij}+O(d|Dw|+1)|u_{ii}|+O(d|Dw|^{3}+|Dw|^{2}),$
Assume that $|Du|$ attains its boundary maximum at $y_{0}\in\partial\Omega$ and
@@
As $\widetilde{G}(x_{0})\geq\widetilde{G}(y_{0})$
(3.15) $\psi(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{1})\text{e}^{-\frac{1}{4}}}{4}|Du|^{2}(y_{0}),$
this follows that
(3.16) $\psi(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{1})}{8C}|Du|^{2}(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{1})}{16C}|Dw|^{2}(x_{0})=C_{0}|Dw|^{2}(x_{0})>0,$
where $C_{0}\in(0,1)$. We choose coordinates such that the matrix $D^{2}u(x_{0})$ is diagonal. It is straightforward to observe that
$\left|\overrightarrow{T}\right|\leq|Dw|,\quad\psi\leq\left|\overrightarrow{T}\right||Dw|,$
where $T_{k}=\sum_{l=1}^{n}C^{kl}w_{l}$ and $\overrightarrow{T}=(T_{1},T_{2},\cdots,T_{n})$. Then
@@
Without loss of generality, we may assume the inequality $T_{1}w_{1}\geq\frac{C_{0}}{n}|Dw|^{2}>0$ holds, which immediately yields $\frac{w_{1}}{T_{1}}\geq\frac{C_{0}}{n}$. Now choose $\mu$ sufficiently small to guarantee that
@@
denote ${C^{kl}}_{,i}=\frac{\partial C^{kl}}{\partial x_{i}}$ and ${C^{kl}}_{,ij}=\frac{\partial^{2}C^{kl}}{\partial x_{i}\partial x_{j}}$, for $i=1$ in equation (3.19), we obtain
(3.20) $\sum_{l=1}^{n}T_{l}w_{l1}=-\frac{\psi}{2}(g^{\prime}u_{1}+\alpha_{1}d_{1})-\sum_{k,l=1}^{n}\frac{{C^{kl}}_{,1}}{2}w_{k}w_{l},$
combine (3.13), (3.17) and (3.20)
$\left(1-\frac{\varphi_{z}d}{\cos\theta}\right)u_{11}=O(d|Du|^{2}+|Du|)\frac{\sum_{k}T_{k}}{T_{1}}-\frac{\psi}{2}\left(g^{\prime}\frac{u_{1}}{T_{1}}+\frac{B_{0}d_{1}}{T_{1}}\right)-\sum_{k,l}\frac{{C^{kl,}}_{1}}{2}\frac{w_{k}w_{l}}{T_{1}},$
it implies that $|u_{11}|\leq C|Du|^{2}$, and we can get $u_{11}<0$ if $\mu$ is small enough. By using the Cauchy-schwarz inequality at $x_{0}$
(3.21) $0\geq F^{ij}\widetilde{G}_{ij}=&\;\frac{F^{ij}\left(\sum_{k,l=1}^{n}C^{kl}w_{k}w_{l}\right)_{ij}}{\psi}-2\alpha_{1}g^{\prime}F^{ij}u_{i}d_{j}-\alpha_{1}^{2}F^{ij}d_{i}d_{j}\\ &\;+\left[g^{\prime\prime}-\left(g^{\prime}\right)^{2}\right]F^{ij}u_{i}u_{j}+g^{\prime}F^{ij}u_{ij}+\alpha_{1}F^{ij}d_{ij}\\ \geq&\;\frac{F^{ij}\left(\sum_{k,l=1}^{n}C^{kl}w_{k}w_{l}\right)_{ij}}{\psi}-2\alpha_{1}^{2}F^{ij}d_{i}d_{j}+\alpha_{1}F^{ij}d_{ij}\\ &\;+\left[g^{\prime\prime}-2\left(g^{\prime}\right)^{2}\right]F^{ij}u_{i}u_{j}+g^{\prime}(k-l)f\\ =&\;\text{I}+\text{II}+\text{III}+\text{IV}+\text{V}.\end{aligned}$
@@
where $K_{0}$ is a positive constant related to the geometry of $\partial\Omega$. It remains to address the complex term I.
@@
For the term $I_{2}$, by (2.2), (3.13), (3.17), Lemma 2.4 and the condition $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$ which is employed to address the lack of $\inf f$, we derive
@@
$\;-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|,$
similarly
@@
(3.29) $\mathcal{H}=2\sum_{i=1}^{n}F^{ii}C^{ii}u_{ii}^{2}-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}|F^{ii}u_{ii}|,$
without loss of generality, we assume that $C^{11}$ be the smallest of $\left\{C^{ii}\right\}_{i=1}^{n}$. This assumption implies that $C^{ii}\geq\frac{1}{2}$ for all $i\geq 2$, otherwise it follows that $\sum_{i=1}^{n}C^{ii}<n-1$, which contradicts with $\sum_{i=1}^{n}C^{ii}=n-1$. A direct computation shows that
(3.30) $\mathcal{H}\geq-C\sum_{i=1}^{n}F^{ii}\left(d|Dw|^{2}+|Dw|\right)^{2},$
thus
@@
Inserting (3.22) and (3.31) into (3.21) yields
(3.32) $0\geq$ $\left\{C_{2}\left[g^{\prime\prime}-2\left(g^{\prime}\right)^{2}\right]|Dw|^{2}-C_{3}\mu|Dw|^{2}-C|Dw|\right\}\sum_{i=1}^{n}F^{ii}$
$\geq$ $\left(\frac{C_{2}}{128M^{2}}|Dw|^{2}-C_{3}\mu|Dw|^{2}-C|Dw|\right)\sum_{i=1}^{n}F^{ii}$
$\geq$ $\left(\frac{C_{2}}{256M^{2}}|Dw|^{2}-C|Dw|\right)\sum_{i=1}^{n}F^{ii},$
where the last inequality follows by taking $\mu$ sufficiently small. This completes the proof. ∎